# Improved Fine-Tuning of Large Multimodal Models for Hateful Meme Detection

### v1

@@ B.2 Hyperparameters for MLP and Stage 2 Fine-tuning
The default hyperparameters for the MLP and the stage 2 RGCL fine-tuning are shown in Table 6. The modeling hyperparameters are based on RGCL’s setting Mei et al. (2024). With this configuration of hyperparameters, the number of trainable parameters is about 5 million.
@@
Appendix C Baseline Methods
@@
  • MOMENTA Pramanick et al. (2021b) leverages trainable fusion layers—such as Cross-Modal Attention Fusion—to integrate multimodal features extracted by CLIP for improved hateful meme detection.
@@
Appendix D Comparing RKC and In-Context Learning under different Fine-tuning Schemes
In this section, we compare the performance of the RKC inference mode against few-shot in-context learning for pre-trained LMMs, SFT LMMs, and LMMs fine-tuned using our proposed LMM-RGCL framework under the cross-dataset setting. As shown in Table 7, RKC consistently outperforms the few-shot in-context learning approach across all LMM variants, indicating that RKC makes better use of demonstration examples. Furthermore, LMM-RGCL fine-tuned LMMs with RKC outperform SFT LMMs with RKC, highlighting the advantages of our fine-tuning strategy.
HatefulMemes HarMeme
Model Mode AUC Acc. AUC Acc.
Pre-trained Few-shot 71.5 63.8 71.5 67.2
Pre-trained RKC 74.5 64.5 80.1 72.4
SFT Few-shot 72.3 60.6 67.2 62.4
SFT RKC 75.8 67.1 84.5 75.4
LMM-RGCL Few-shot 70.8 63.5 73.2 68.1
LMM-RGCL RKC 77.1 69.3 88.8 81.7
Table 7: Comparing Pre-trained, SFT and LMM-RGCL systems with few-shot learning and RKC with Qwen2VL-7B under cross-dataset settings. See Table 2
Appendix E Ablation study on the loss function
HatefulMemes HarMeme
Mode AUC Acc. AUC Acc.
LMM-RGCL 91.1 82.1 93.2 88.1
w/o $\mathcal{L}^{LM}$ in stage 1 88.4 79.6 90.9 85.1
w/o $\mathcal{L}^{RGCLL}$ in stage 2 90.2 81.2 91.9 86.4
w/o $\mathcal{L}^{LR}$ in stage 2 89.2 80.6 91.6 87.2
(a) Supervised settings, see Table 1
HatefulMemes HarMeme
Mode AUC Acc. AUC Acc.
LMM-RGCL 77.1 69.3 88.8 81.7
w/o $\mathcal{L}^{LM}$ in stage 1 75.4 66.6 87.3 81.1
w/o $\mathcal{L}^{RGCLL}$ in stage 2 73.8 64.3 82.9 76.5
w/o $\mathcal{L}^{LR}$ in stage 2 76.4 67.9 86.9 80.6
(b) Cross-dataset settings, see Table 2
Table 8: Ablation study of LMM-RGCL two-stage fine-tuning framework on Qwen2VL-7B, evaluating the impact of removing any of the loss objectives.
Appendix F Case Analysis
Case 1 Case 2 Case 3
Meme [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Ground Truth #Hateful #Hateful #Hateful
SFT #Benign #Benign #Benign
LMM-RGCL #Hateful #Hateful #Hateful
Case 4 Case 5 Case 6
Meme [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Ground Truth #Hateful #Benign #Benign
SFT #Benign #Hateful #Hateful
LMM-RGCL #Hateful #Benign #Benign
Table 9: Visualization of cases from SFT Qwen2VL-7B and LMM-RGCL Qwen2VL-7B Models on the HatefulMemes Dataset. Case 5 contains an insect in the meme; we applied a blurring filter to obscure it.
Case 1 Case 2 Case 3
Meme [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Ground Truth #Hateful #Hateful #Hateful
SFT #Benign #Benign #Benign
LMM-RGCL #Benign #Benign #Benign
Table 10: The error cases of SFT Qwen2VL-7B and LMM-RGCL Qwen2VL-7B models on HatefulMemes dataset
F.1 Comparing SFT and LMM-RGCL Predictions
Table 9 presents examples where our LMM-RGCL method successfully corrects prediction errors made by the SFT model on Qwen2VL-7B. Cases 1-4 involve hateful memes, while Cases 5-6 are benign memes that the SFT model misclassified, primarily due to poor multimodal alignment. These examples require a deep, joint understanding of both the image and text, a challenge that our LMM-RGCL effectively addresses. For example, in Case 2, the model needs to use its understanding of Japanese culture and associate this knowledge with the visual cues in the image.
F.2 Error Analysis
In Table 10, we present examples where LMM-RGCL was unable to correct errors made by the baseline SFT model. In the first case, the model struggles with the nuanced visual understanding required to interpret the disabled body of the swimmer. Additionally, these examples demand complex reasoning to assess the hatefulness of the memes. Interpreting such nuanced meanings remains a challenge for current models. However, we anticipate that the advanced reasoning capabilities of emerging systems like OpenAI-o1 OpenAI (2024b) and DeepSeek-R1 DeepSeek-AI (2025) will help address these limitations.
Appendix G AI Assistance

### v2

@@ B.2 Hyperparameters for MLP and Stage 2 Fine-tuning
The default hyperparameters for the MLP and the stage 2 contrastive fine-tuning are shown in Table 6. With this configuration of hyperparameters, the number of trainable parameters is about 5 million.
@@
Appendix C Statistical Significance Test
We conduct statistical significance tests comparing the performance of SFT and RA-HMD fine-tuned LMMs, as reported in Table 7.
HatefulMemes HarMeme MAMI Harm-P MultiOFF PrideMM
Model AUC Acc. AUC Acc. AUC Acc. Acc. F1 Acc. F1 Acc. F1
LLaVA-1.5-7B
[…]
Qwen2VL-2B
[…]
Qwen2VL-7B
[…]
Table 7: For each LMM, we provide the  p-value from significance testing between SFT and RA-HMD.
Appendix D Insights for using LMMs representation for Meme Classification
There has been substantial interest in adapting decoder-only language models for retrieval and classification tasks. In this section, we summarize the novelty of our approach in comparison to previous efforts that attempt to repurpose decoder-only large language models (LLMs) or large multimodal models (LMMs) for such tasks.
We categorize our adaptation attempts into two groups: (1) those that entirely failed to work for hateful meme classification, and (2) those that showed some promise in improving classification performance but significantly compromised the model’s language generation capabilities. Finally, we explain the rationale that led to the design of our current RA-HMD architecture.
Attempts That Do Not Work
Some prior approaches aim to modify the model architecture with contrastive training objective to enable bidirectional representations from decoder-only models.
For instance, LLM2Vec BehnamGhader et al. (2024) proposes to introduce bidirectional attention and masked next-token prediction to make decoder-only models suitable for retrieval and general-purpose text embeddings.
AutoVER Xiao et al. (2024), on the other hand, introduces a special [RET] token and learns its embedding for visual entity retrieval.
However, neither approach yielded meaningful improvements in our setting. When applied to hateful meme classification, both methods achieved classification accuracies only slightly above 60%. We attribute this failure to the limited availability of hateful meme training data, which is insufficient for training these complex adaptations effectively.
Approaches That Partially Work
Other works have explored using decoder-only LLM or LMM embeddings by either pooling the output token representations or using the final-layer embedding of the last token Zhang et al. (2024); Li et al. (2023b); Wang et al. (2024a), often in combination with contrastive learning.
In our experiments, we found that mean pooling generally underperformed compared to using the final token embedding. While this approach achieved a reasonable classification accuracy of about 77% on the HatefulMemes dataset using Qwen2VL-7B (RA-HMS Qwen2VL-7B has an accuracy of 82%), it had a major drawback: the language generation ability of the model was completely compromised. In practice, the model became unable to generate coherent text, suggesting that the learned representations were no longer aligned with the original language modeling objective.
Our approach
To address these limitations, we propose two key improvements that underpin the RA-HMD architecture:
  • MLP Projection Head: We introduce a lightweight, trainable multilayer perceptron (MLP) on top of the final token embedding from the LMM. While the language model head (LMH) continues to use the original last-token embedding for language generation, the classification and retrieval heads operate on the MLP-projected embedding. This separation enables the model to retain its language generation capability while learning representations that are better suited for classification and retrieval.
  • Preserving the Language Modeling Objective: During Stage 1 of RA-HMD fine-tuning, we retain the original language modeling loss in the training objective. This encourages the base embeddings to remain useful for text generation, avoiding overfitting solely to the classification task and preserving the model’s general-purpose functionality.
Appendix E Numbers of Shots and Neighbors
Figure 3 shows that as the number of nearest neighbors $K$ for RKC increases, the performance continues to increase for both AUC and accuracy, plateauing at around $K=20$. The consistent improvement in performance indicates that RKC trained with RA-HMD utilizes demonstration examples more effectively than the standard in-context learning framework.
Figure 3: Effects of increasing number of shots for in-context learning with pre-trained LMM and effects of increasing top K nearest neighbors for RKC trained with RA-HMD
Appendix F Comparing Out-of-Domain Generalization Across Model Variants with In-Context Learning and RKC
In this section, we compare the performance of the RKC inference mode against few-shot in-context learning for pre-trained LMMs, SFT LMMs, and LMMs fine-tuned using our proposed RA-HMD framework under the cross-dataset setting in Table 8. We observe three things:
  • RA-HMD-trained Qwen2VL-7B exhibits more robust generalization in the cross-dataset setting. Its in-context learning performance in few-shot scenarios consistently surpasses that of other models.
  • RKC consistently outperforms the in-context learning approach across all LMM variants, demonstrating its superior ability to leverage demonstration examples.
  • Moreover, RA-HMD fine-tuned LMMs with RKC outperform SFT LMMs with RKC, highlighting the effectiveness of our fine-tuning strategy.
HatefulMemes HarMeme
Model Mode AUC Acc. AUC Acc.
Pre-trained Few-shot 71.5 63.8 71.5 67.2
Pre-trained RKC 74.5 64.5 80.1 72.4
SFT Few-shot 72.3 60.6 67.2 62.4
SFT RKC 75.8 67.1 84.5 75.4
RA-HMD Few-shot 74.3 63.5 73.2 68.1
RA-HMD RKC 77.1 69.3 88.8 81.7
Table 8: Comparing Pre-trained, SFT and RA-HMD systems with few-shot learning and RKC with Qwen2VL-7B under cross-dataset settings. See Table 2
Appendix G Comparing Different Inference Modes
Table 9 compares Qwen2VL-7B fine-tuned with RA-HMD using the three classifiers. Our results indicate that under supervised settings, the differences among the three inference modes are minimal. However, under cross-dataset settings, there is a significant disparity in generalization performance. Notably, RKC outperforms both LMH and LRC, underscoring its superior effectiveness in handling out-of-domain examples.
HatefulMemes HarMeme
Inference Mode AUC Acc. AUC Acc.
LMH 90.2 81.9 92.8 88.0
LRC 91.1 82.1 93.2 88.1
RKC 90.8 81.8 93.2 88.0
(a) Supervised settings, see Table 1 for detailed settings
HatefulMemes HarMeme
Inference Mode AUC Acc. AUC Acc.
LMH 74.2 64.3 64.5 60.3
LRC 59.5 55.4 57.9 52.2
RKC 77.1 69.3 88.8 81.7
(b) Cross-dataset settings, see Table 2 for detailed settings
Table 9: Comparing different inference modes using RA-HMD fine-tuned Qwen2VL-7B. RKC shows much better out-of-domain generalization compared to other inference modes.
Appendix H Ablation study on the loss function
Furthermore, Mei et al. (2024) utilize in-batch negative examples alongside hard negative examples during training. However, we find that incorporating in-batch negatives in Stage 2 of RA-HMD’s contrastive fine-tuning introduces noise and leads to a slight degradation in performance.
HatefulMemes HarMeme
Mode AUC Acc. AUC Acc.
RA-HMD 91.1 82.1 93.2 88.1
w/o $\mathcal{L}^{LM}$ in stage 1 88.4 79.6 90.9 85.1
w/o $\mathcal{L}^{RGCL}$ in stage 2 90.2 81.2 91.9 86.4
w/o $\mathcal{L}^{LR}$ in stage 2 89.2 80.6 91.6 87.2
(a) Supervised settings, see Table 1
HatefulMemes HarMeme
Mode AUC Acc. AUC Acc.
RA-HMD 77.1 69.3 88.8 81.7
w/o $\mathcal{L}^{LM}$ in stage 1 75.4 66.6 87.3 81.1
w/o $\mathcal{L}^{RGCL}$ in stage 2 73.8 64.3 82.9 76.5
w/o $\mathcal{L}^{LR}$ in stage 2 76.4 67.9 86.9 80.6
(b) Cross-dataset settings, see Table 2
Table 10: Ablation study of RA-HMD two-stage fine-tuning framework on Qwen2VL-7B, evaluating the impact of removing any of the loss objectives.
Appendix I Evaluation on General Vision-Language Benchmarks
To evaluate general vision-language capabilities, we use models fine-tuned on the HatefulMemes dataset via both supervised fine-tuning (SFT) and our proposed RA-HMD approach. We conduct the evaluation using the VLMEvalKit package [^5] https://github.com/open-compass/VLMEvalKit Duan et al. (2024). For MMMU, we report accuracy based on exact match using the MMMU_DEV_VAL split to ensure reproducibility. For SeedBench, we use the SeedBench_IMG subset, also evaluated with exact match. For GQA, we report results on the GQA_TestDev_Balanced split. The full results are provided in Table 11.
Model MMMU SEEDBench GQA
Qwen2VL-2B 40.2 72.7 60.4
+SFT 39.1 72.1 57.0
+RA-HMD 40.4 72.7 60.1
Qwen2VL-7B 49.3 76.4 62.4
+SFT 48.0 75.2 61.2
+RA-HMD 49.2 76.4 62.3
Table 11: Comparison of the pretrained, SFT, and RA-HMD Qwen2VL-2B models on three general vision-language benchmarks. The SFT and RA-HMD models are fine-tuned on the HatefulMemes dataset.
Appendix J GPT-4o Results and Prompts
We adopt the gpt-4o-2024-05-13 model for our evaluations. We explore two prompting strategies for GPT-4o. In Table 1, we present results using optimized prompts, where the specific type of hate (e.g., hatefulness, harmfulness, offensiveness, misogyny, etc.) is explicitly provided based on the characteristics of each dataset. Below, we include the detailed prompt formulations used for each dataset:
  • HatefulMemes:
    You are an expert in identifying meme-based social abuse. Given the meme, with the Text [T] accompanied by the image. The task is to classify whether the meme contains hateful content or not. Please answer with ’yes’ or ’no’.
  • Harmeme:
    You are an expert in identifying meme-based social abuse. Given the meme, with the Text [T] accompanied by the image. The task is to classify whether the meme contains harmful content or not. Please answer with ’yes’ or ’no’.
  • Harmp:
    You are an expert in identifying meme-based social abuse. Given the meme, with the Text [T] accompanied by the image. The task is to classify whether the meme contains harmful content or not. Please answer with ’yes’ or ’no’.
  • MultiOFF:
    You are an expert in identifying meme-based social abuse. Given the meme, with the Text [T] accompanied by the image. The task is to classify whether the meme contains offensive content or not. Please answer with ’yes’ or ’no’.
  • PrideMM:
    You are an expert in identifying meme-based social abuse. Given the meme, with the Text [T] accompanied by the image. The task is to classify whether the meme contains hateful content related to LGBTQ+ Pride movement or not. Please answer with ’yes’ or ’no’.
  • MAMI:
    You are an expert in identifying meme-based social abuse. Given the meme, with the Text [T] accompanied by the image. The task is to classify whether the meme contains misogyny or not. Please answer with ’yes’ or ’no’.
The low-resource comparison in Table 2 is designed to reflect real-world scenarios to detect the evolving harmful memes on the internet, where the specific type of hate is often unknown. Accordingly, we use the general term “hate” across all six datasets in this setting. Below, we include the detailed prompt formulations:
  • Given the meme, with the Text [T] accompanied by the image. Does the meme contain any hateful content or any social abuse?
We directly present the comparison between the two sets of results in Table 12. For reference, Goat-Bench Lin et al. (2025) published GPT-4o results on similar datasets using task-specific prompts. However, their evaluation is based on different data splits, making the results not directly comparable. Below, we summarize key differences in performance:
  • Hatefulness: This benchmark corresponds to our HatefulMemes dataset. While we use the test_seen split, Goat-Bench uses the test_unseen split. Despite this difference, the results are comparable: they report an accuracy of 71.7, while ours is 71.3.
  • Harmfulness: This benchmark aligns with our combined evaluation on HarMeme and Harm-P. Goat-Bench reports an accuracy of 66.01, whereas we achieve 72.9 on HarMeme and 63.1 on Harm-P.
  • Offensiveness: Due to the small size of the MultiOFF dataset, Goat-Bench evaluates performance using the combined training, validation, and test sets. In contrast, we report results based solely on the test set. They report an accuracy of 62.13 and an F1 score of 61.16, compared to our accuracy of 58.3 and F1 score of 58.1.
HatefulMemes HarMeme MAMI Harm-P MultiOFF PrideMM
Model AUC Acc. AUC Acc. AUC Acc. Acc. F1 Acc. F1 Acc. F1
GPT-4o (Specific Prompt) - $71.3$ $72.9$ $79.4$ $63.1$ $64.5$ $58.3$ $58.1$ $75.3$ $73.7$
GPT-4o (General Prompt) - $66.40$ $68.4$ $72.90$ $55.39$ $55.11$ $61.07$ $51.07$ $63.79$ $62.28$
Table 12: Comparing the performance with GPT-4o with different prompts
Appendix K Rationale Generated by LMMs
We compare meme explanations generated by the SFT and RA-HMD fine-tuned Qwen2VL-7B models (Table 1 rows 15 and 16). RA-HMD is evaluated after Stage-1 fine-tuning, as Stage-2 does not fine-tune the Language Model Head for language generation. For evaluation, we prompt the above two models with:
"Does the meme contain any hate speech or offensive content? Please provide a detailed explanation."
Following prior work Yang et al. (2023), we assess explanation quality using an LLM judge. Specifically, we provide GPT-4o-mini (gpt-4o-mini-2024-07-18) with reference explanations from Hee et al. (2023) and use the comparison template from Yang et al. (2023) for pairwise evaluation. For reference, the prompt we use is:
[Instruction]
Please act as an impartial judge and evaluate the quality of the responses provided by two AI assistants to the user question displayed below. You should choose the assistant that follows the users instructions and answers the users question more accurately.
When choosing the assistant, please consider the true answers below:
Target: {true_label}{reference_info}
Implied Statement: This is a hateful meme with text: "{meme_text}"
Your evaluation should consider which response is more similar to the true answers. Begin your evaluation by comparing the two responses and provide a short explanation. Avoid any positional biases and ensure that the order in which the responses were presented does not influence your decision. Do not allow the length of the responses to influence your evaluation. Do not favor certain names of the assistants. Be as objective as possible. After providing your explanation, output your final verdict by strictly following this format: "[[A]]" if assistant A is more accurate, "[[B]]" if assistant B is more accurate, and "[[C]]" for a tie.
[Question]
Determine whether the following post is offensive or not. You must choose one of the options.
Post: A meme with text: "{meme_text}"
Options:
(A) Offensive
(B) Not offensive
[The Start of Assistant As Answer]
{example_a["reasoning"]}
[The End of Assistant As Answer]
[The Start of Assistant Bs Answer]
{example_b["reasoning"]}
[The End of Assistant Bs Answer]}
Results are as follows:
  • RA-HMD beats SFT: 61.5
  • RA-HMD ties SFT: 13.8
  • SFT beats RA-HMD: 24.7
Based on the analysis of the generated explanations, we find that improvements in classification accuracy are supported by deeper semantic understanding of memes. For challenging examples, where comprehension of background events or fine-grained details from the image is required, the RA-HMD fine-tuned system generates better explanations.
For reference, we include sample explanations of the meme examples in the Case Analysis in Appendix L Table 13.
Case 1 (Meme references 9/11 in a political context)
  • SFT: No. It does not contain any hate speech or offensive content. It appears to be a political statement questioning the decision to vote for a particular political figure or party. It suggests a critique of the decision to support a political figure or party that is associated with the events mentioned in the text.
  • RA-HMD: Yes. The meme contains hate speech. It is offensive and disrespectful to suggest that the Twin Towers should be voted into the government. It is a tragic event that should never be forgotten, and it is not appropriate to use it as a political meme.
Case 2 (Meme associates modern Japanese people with suicide)
  • SFT : No. It does not contain any hate speech or offensive content. It simply features a person holding a gun to their head with the text "modern japanese" and two quotation marks. The text and imagery do not convey any offensive or hateful messages.
  • RA-HMD: Yes. The meme contains hate speech. It is a hateful meme that associates modern Japanese people with suicide. It is a very offensive meme.
Case 3 (Meme objectifies a woman with references to the disability)
  • SFT: No. It does not contain any hate speech or offensive content. The meme contains text that implies a sexual service is available for purchase, which could be interpreted as objectifying or degrading. However, the image itself does not contain any explicit or offensive content.
  • RA-HMD: Yes. The meme contains hate speech. It is a hateful meme that makes fun of a woman’s disability. It is a very offensive meme.
These results further demonstrate the superior meme understanding capabilities of RA-HMD, especially when compared to SFT.
Case 1 Case 2 Case 3
Meme [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Ground Truth #Hateful #Hateful #Hateful
SFT #Benign #Benign #Benign
RA-HMD #Hateful #Hateful #Hateful
Case 4 Case 5 Case 6
Meme [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Ground Truth #Hateful #Benign #Benign
SFT #Benign #Hateful #Hateful
RA-HMD #Hateful #Benign #Benign
Table 13: Visualization of cases from SFT Qwen2VL-7B and RA-HMD Qwen2VL-7B Models on the HatefulMemes Dataset. Case 5 contains an insect in the meme; we applied a blurring filter to obscure it.
Case 1 Case 2 Case 3
Meme [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Ground Truth #Hateful #Hateful #Hateful
SFT #Benign #Benign #Benign
RA-HMD #Benign #Benign #Benign
Table 14: The error cases of SFT Qwen2VL-7B and RA-HMD Qwen2VL-7B models on HatefulMemes dataset
Appendix L Case Analysis
L.1 Comparing SFT and RA-HMD Predictions
Table 13 presents examples where our RA-HMD method successfully corrects prediction errors made by the SFT model on Qwen2VL-7B. Cases 1-4 involve hateful memes, while Cases 5-6 are benign memes that the SFT model misclassified, primarily due to poor multimodal alignment. These examples require a deep, joint understanding of both the image and text, a challenge that our RA-HMD effectively addresses. For example, in Case 2, the model needs to use its understanding of Japanese culture and associate this knowledge with the visual cues in the image.
L.2 Error Analysis
In Table 14, we present examples where RA-HMD was unable to correct errors made by the baseline SFT model. In the first case, the model struggles with the nuanced visual understanding required to interpret the disabled body of the swimmer. Additionally, these examples demand complex reasoning to assess the hatefulness of the memes. Interpreting such nuanced meanings remains a challenge for current models. However, we anticipate that the advanced reasoning capabilities of emerging systems like OpenAI-o1 OpenAI (2024b) and DeepSeek-R1 DeepSeek-AI (2025) will help address these limitations.
Appendix M Baseline Methods
@@ Appendix I Evaluation on General Vision-Language Benchmarks
Appendix N AI Assistance